True or False
The rule y^2 = x^3 defines a function.

Answer by Theo(13342)

i would say false.

y^2 = x^3 is solved as follows:

take the square root of both sides to get:

y = plus or minus sqrt(x^3)

that plus or minus tell you that you have more than one value of y for each value of x.

for example:

assume that x = 4.

you have y^2 = x^3 which means that y^2 = 64 when x is 4 because 4^3 = 64.

take the square root of both sides of that equation to get y = +/- 8.

for the value of x = 4, you can get y = + 8 or - 8.

more than one value of y for a given value of x means that you do not have a function since that violates the rules of a function that state that there is one and only one value of y for each and every value of x.

there may be only one possible value of y^2, but there are more than one possible value of y.